rD= 55% βE=.10 risk free =4% r m=19% D/V=0.6 beta of debt?
Answer #1

Usually debt beta id assumed to be zero when calculating WACC>

But ideally, taking debt beta as 0 means that it has no market risk. Which is usually not true.

Debt beta can be calculated using CAPM, as CAPM can be used to price any asset, so if we are given the cost of debt, we can calculate the debt beta.

Thus, debt beta= (Kd-Rf)/RM, where Kd=cost of debt, Rf=risk free rate, Rm=Market risk premium

(55%-4%)/19%=2.68

Please reach out for any clarifications